# Self-lensing of moving gravitational-wave sources can break the microlensing crossing timescale degeneracy

### v1

@@ Abstract
When a moving gravitational-wave (GW) source travels behind a massive astrophysical object, its signal is gravitationally lensed, showing a waveform distortion similar to a Paczyński curve. We present a first study of the lensing signature of a massive black hole (MBH) on a frequency-dependent GW signal from a moving binary merger. For both light and GW sources in a Keplerian circular orbit around a MBH lens, the self-lensing geometry breaks the microlensing degeneracy in the Einstein radius crossing timescale $t_{\rm E}$. The duration of the curve ($2t_{\rm E}$) becomes independent on the MBH mass $M_{\rm MBH}$, and provides a direct measure of the distance $d_{\rm LS}$ to the MBH. However, $M_{\rm MBH}$ remains unknown. We show that, in GW signals, the redshifted mass $M_{{\rm MBH},z}$ can additionally be obtained from the interference pattern, by measuring the modulation period $T$, the GW frequency $f$, and $t_{\rm E}$: $M_{{\rm MBH},z}\simeq 2.5\times 10^{6}\,M_{\odot}\,(t_{\rm E}/[100\,{\rm s}])\,(f\,T)^{-1}$. If this lensing signature is not considered, it may be confused with other waveform distortions, especially in the modeling of overlapping signals in next generation ground-based GW detectors. The observation of one of these curves and its associated parameters may help (1) constrain the orbital distance $d_{\rm LS}$ of sources, especially around low mass MBHs at the centers of star clusters and galaxies, (2) additionally estimate the mass $M_{{\rm MBH},z}$ of these MBHs, and (3) infer the orbital inclination of the binary. Simultaneously obtaining $d_{\rm LS}$ and $M_{{\rm MBH},z}$ through self-lensing can help constrain the astrophysical environments where GW signals come from.
@@ I Introduction
When light and gravitational waves (GWs) travel through the space-time curvature of massive astrophysical objects, they are deflected and distorted by gravitational lensing. Lensing effects have been detected with light [walsh-79, alcock-93, udalski-93], although they have not confidently been detected with GWs yet [hannuksela-19, dai-20, LVK-O3a-lensing, LVK-O3ab-Lensing, janquart-23] (though see [231123]). GWs from binary mergers, usually refered to as compact binary coalescences (CBCs), have been detected by ground based detectors from the LIGO-Virgo-KAGRA (LVK) Collaboration. To date, nearly $300$ GW events have been confidently detected [LIGO18-GWTC1, LIGO20-GWTC2, LIGO21-GWTC3, venumadhav-20, zackay-21, LIGO25-GWTC4]. As the number of GW events from CBCs increases, the lensing effect is expected to be detected on GWs soon [ng-18, li-18, oguri-18, xu-ezquiaga-holz-22, mukherjee-21, wierda-21].
The approaches to the observation of lensing of light and lensing of GWs are generally different, due to the nature of the detection methods. For light, lensing effects appear as either (1) multiple images of the same source, resolved through imaging (strong lensing); (2) transient increase of the source brightness following a Paczyński curve [paczynski-86], when the source temporarily travels behind the lens, seen through photometry (microlensing); (3) astrometric microlensing. For transient GWs detectable in ground-based detectors, the equivalent of strong lensing is the detection of multiple images of a static source in time domain. Given the transient nature of the detected compact binary merger (CBC) sources, the equivalent to microlensing by low mass lenses is probed through wave optics distortions for static sources [nakamura-98, takahashi-03], rather than moving sources. The equivalent of the Paczyński curve is usually not considered for transient GWs from CBCs in ground-based detectors. Studies either consider a static source and varying frequency [e.g., liao-19, sun-fan-19, hou-20, biesiada-harikumar-21] or a constant frequency and a moving source, the latter of which has been analyzed for continuous GW sources [e.g., depaolis-01, guo-24, sudhagar-25, yang-25, li-25]. Joint studies have been done for space-based detectors [dorazio-loeb-20, yu-21]. To the best of the author’s knowledge, no previous studies have been done for a frequency-dependent moving source in ground-based detectors.
@@
The observation of the lensing curve can give us information of the parameters of the lensing system. In microlensing of light, the duration of the curve (twice the Einstein crossing timescale $t_{\rm E}$) usually has a degeneracy between the lens mass, a combination of the distances, and the relative velocity of the lens and the source. The degeneracy may be broken by measuring the parallax [refsdal-66, gould-94b, gould-00], simultaneous parameters [lee-17, gould-19, for a review], finite-source effects [gould-94a, nemiroff-94], or relativistic effects [rahvar-20]. By contrast, when the source is orbiting a massive black hole (MBH) that acts as the lens (self-lensing), the Einstein crossing timescale is only dependent on the orbital distance $d_{\rm LS}$ from the source to the lens [rahvar-11]. For both light and GWs, the breaking of the Einstein crossing timescale degeneracy thus provides an independent way to measure the orbital distance $d_{\rm LS}$ of the source.
@@ II Lensing curve on a chirping signal
Figure 1: Schema of the lensing configuration considered in this study. Top: The GW source (S) is a compact binary orbiting a MBH at a distance $d_{\rm LS}$. The MBH acts as a gravitational lens (L), affecting the GW signal, with arrives distorted at the observer (O) when S, L and O are close to alignment. Bottom: View from the observer, where the source travels behind the lens. The lensing effect is dominant behind the cross-section of radius $R_{\rm E}$ and, as we will see in Sec. IV, the amplitude of the signal is modulated by the transmission factor ($F$ in Eq. (3)) shown by the pattern here. The source can have an offset $y_{0}R_{\rm E}$ with respect to the perfect alignment (dashed line going through the center of the lens).
The source travels behind the lensing cross-section, generally [^1] Following classical microlensing conventions, we are taking a maximum source position of $y_{\rm max}=1$ as the cross-section threshold. For GWs, the detectability thresholds for $y$ can be different (usually $\gtrsim 1$) than for electromagnetic lensing, and in that case the cross-section is $\pi R_{\rm E}^{2}y_{\rm max}^{2}$. given by $\pi R_{\rm E}^{2}$, where $R_{\rm E}$ is the Einstein radius, which for self-lensing is simplified to $R_{\rm E}\simeq\sqrt{2R_{\rm S}d_{\rm LS}}$  [^2] Throughout this work, due to self-lensing, the distance $d_{\rm L}$ from the observer to the lens can be taken as practically equal to the distance $d_{\rm S}$ from the observer to the source., with $R_{\rm S}=2GM_{\rm MBH}/c^{2}$ being the Schwarzschild radius of the MBH of mass $M_{\rm MBH}$. $G$ is the gravitational constant and $c$ is the speed of light in vacuum. The source is generally not perfectly aligned (i.e., not crossing exactly behind the lens), it has a physical offset $\eta_{0}\simeq y_{0}R_{\rm E}$ where $y_{0}$ is the dimensionless source position (“impact parameter”, sometimes known as $u_{0}$) of closest approach to the lens at time $t_{0}$. As the source moves, the source position $y$ with respect to the lens changes over time ($t$),
@@ II.1 The crossing timescale depends directly on the orbital distance
Compared to a general source, this result is not degenerate and provides a direct measure of the distance $d_{\rm LS}$, as we analyze in the following.
@@ II.2 Lensing signature on the GW waveform
A lensed CBC signal is shown in Fig. 2. To interpret it, we start chronologically with the source far from alignment, when the first image looks unlensed ($\sqrt{\mu_{1}}\rightarrow 1$), the second image is very faint ($\sqrt{\mu_{2}}\rightarrow 0$) and the time delay between them is large. As the source moves behind the lens, both images are magnified and the time delay becomes shorter. At $t_{0}$, the images arrive simultaneously and at their maximum joint amplification, corresponding to the peak. As the source moves away, the situation is reversed back to the same as in the beginning, in this case the first image gets demagnified until it is too faint ($\sqrt{\mu_{1}}\rightarrow 0$) and the second image looks unlensed, $\sqrt{\mu_{2}}\rightarrow 1$.
@@
Figure 2: Representation of a GW signal CBC “chirp”, through the strain $h$ (relative deformation of space-time) as a function of time $t$. The amplitude of the Paczyński-like curve on the inspiral phase of the signal can be comparable (or even larger) than the final merger. The parameters used here are a source “chirp mass” $M_{\rm chirp}=1\,{\rm M}_{\odot}$, a distance from the observer $d_{\rm L}=1\,{\rm Gpc}$ (just affecting the overall amplitude of the strain), $y_{0}=0.1$ and $d_{\rm LS}\simeq 3\times 10^{10}\,{\rm m}$. In self-lensing, the width of the lensing curve ($2t_{\rm E}=4d_{\rm LS}/c$) is independent on $M_{\rm MBH}$ [Eq. (2)]. Instead, as we will see in Sec. IV, the signal is modulated by an interference pattern, which does depend on $M_{\rm MBH}$ (Fig. 3). For comparison purposes, the lensed signal is represented in the time reference of the arrival time of the first image, $t_{1}$. For simplicity, we represent the inspiral part of the signal from $f\gtrsim 12\,{\rm Hz}$, truncated at merger.
@@
Although the source waveform $h_{\rm UL}(t)$ is chosen to be a CBC “chirp”, with frequency dependence $f(t)\propto M_{\rm chirp}^{-5/8},$ the method can be applied to any smoothly varying source. In Fig. 2, the oscillations are too narrow to distinguish the interference pattern or the phase. A zoom view is seen in Fig. 3: while the interference pattern modulations are distinguished, the oscillations still have too many cycles to be resolved by eye. In this work, the amplitude already gives us enough information. However, the phase may give additional information about environmental phase shifts, which are not taken into account here and may be required in more exhaustive searches.
@@
Figure 3: Zoom-in view of Fig. 2, at the peak of the lensing curve at GW frequencies $f\simeq 15\,{\rm Hz}$. The interference pattern modulating the GW signal depends on the mass, (a) $M_{{\rm MBH},z}=10^{5}\,M_{\odot}$, (b) $M_{{\rm MBH},z}=10^{6}\,M_{\odot}$ and the distance $d_{\rm LS}\simeq 3\times 10^{10}\,{\rm m}$ (or equivalently $t_{\rm E}\simeq 200\,{\rm s}$). The modulation period far from the peak can be estimated with Eq. (15), for example here in the segment $t\in(400-420)\,{\rm s}$.
@@
In ground-based GW detectors, this lensing curve on the GW signal may appear as an excess power. If not taken into account, the full signal may be missed or misinterpreted. Although in most cases it might not be confused with short bursts (glitches or other burst signals such as eccentric pulses/bursts [samsing-25-GWpulsar]) because $t_{\rm E}$ is generally longer than their duration, the lensing curve should be taken into account when treating multiple overlapping signals in next-generation detectors.
@@ III Obtaining the orbital distance
Observability for sources of light. The light-curve may be seen for long timescales, provided that the source is at a distance $d_{\rm LS}$ larger than the tidal disruption radius of the source $R_{\rm tidal}$. For typical $R_{\rm tidal}$ of stars around MBHs, we can expect $t_{\rm E}\gtrsim 3\,{\rm hours}$. Low mass MBHs may be favored due to their potential to produce short timescales because $d_{\rm LS}>R_{\rm tidal}\propto M_{\rm MBH}^{1/3}$, which can be interesting for the MBH at the center of our Galaxy (expanded in the Discussion, VI) or intermediate-mass black holes (IMBHs) at the centers of star clusters.
@@
Such close distances may be reached in active galactic nuclei (AGN) disks, if the sources are in migration traps [mckernan-12, bellovary-16, secunda-19]. For example, fixing the distance at $d_{\rm LS}\sim 10\,R_{\rm S}$ [peng-21, leong-25], for $M_{\rm chirp}=1\,M_{\odot}$, a complete curve appears for $M_{\rm MBH}\lesssim 3\cdot 10^{6}\,M_{\odot}$.
@@ IV Obtaining the MBH mass
The lensed GW signal presents a modulation of the amplitude which can be seen in Fig. 3 (a zoom-in view of Fig. 2). The modulation is caused by the phase difference between the lensed images, which creates an interference pattern that can also be found in continuous GW signals [sudhagar-25]. The period of this modulation can be roughly estimated in the GO approximation from Eq. (7), $T=t/(f\Delta t)$ [savastano-24], using $y\sim t/t_{\rm E}$. This period is only formally valid for the radial spacing (interference pattern in Fig. 1). Since in our case the source has an offset $y_{0}$, the period will asymptotically match the modulations far from the $t_{0}$ peak. Nevertheless, the point of this estimation is to show that the period is inversely proportional to the mass of the MBH,
@@
By observing the modulation period $T$ and the GW frequency $f$  [^3] The frequency is assumed to be directly measurable by the detector from the incoming GWs., we can obtain the mass of the MBH:
@@
For the astrophysical parameters we consider (typical MBH masses and LVK/ET/CE frequencies), these modulations are expected to last $\gtrsim{\rm s}$. The interference pattern is thus potentially resolvable in time domain in current ground-based GW detectors.

### v2

@@ Abstract
When a moving gravitational-wave (GW) source travels behind a massive astrophysical object, its signal is gravitationally lensed, showing a waveform distortion similar to a Paczyński curve. We present a first study of the lensing signature of a massive black hole (MBH) on a frequency-dependent GW signal from a moving binary merger. For both light and GW sources in a Keplerian circular orbit around a MBH lens, the self-lensing geometry breaks the microlensing degeneracy in the Einstein radius crossing timescale $t_{\rm E}$. The duration of the curve ($2t_{\rm E}$) becomes independent on the MBH mass $M_{\rm MBH}$, and provides a direct measure of the distance $d_{\rm LS}$ to the MBH. However, $M_{\rm MBH}$ remains unknown. We show that, in GW signals, the redshifted mass $M_{{\rm MBH},z}$ can additionally be obtained from the interference pattern, by measuring the modulation period $T$, the GW frequency $f$, and $t_{\rm E}$: $M_{{\rm MBH},z}\simeq 2.5\times 10^{6}\,M_{\odot}\,(t_{\rm E}/[100\,{\rm s}])\,(f\,T)^{-1}$. If this lensing signature is not considered, it may be confused with other waveform distortions, especially in the modeling of overlapping signals in next generation ground-based GW detectors. The observation of one of these curves and its associated parameters may help (1) constrain the orbital distance $d_{\rm LS}$ of sources, especially around low-mass MBHs at the center of star clusters and galaxies, (2) additionally estimate the mass $M_{{\rm MBH},z}$ of these MBHs, and (3) infer the orbital inclination of the binary. Simultaneously obtaining $d_{\rm LS}$ and $M_{{\rm MBH},z}$ through self-lensing can help constrain the astrophysical environments where GW signals come from.
@@ I Introduction
When light and gravitational waves (GWs) travel through the space-time curvature of massive astrophysical objects, they are deflected and distorted by gravitational lensing. Lensing effects have already been detected on light [walsh-79, alcock-93, udalski-93], although they have not been confidently detected on GWs yet [hannuksela-19, dai-20, LVK-O3a-lensing, LVK-O3ab-Lensing, janquart-23] (though see [231123]). GWs from binary mergers, usually refered to as compact binary coalescences (CBCs), have been detected by ground-based detectors from the LIGO-Virgo-KAGRA (LVK) Collaboration. To date, more than $300$ GW event candidates have been detected [LIGO18-GWTC1, LIGO20-GWTC2, LIGO21-GWTC3, venumadhav-20, zackay-21, LIGO25-GWTC4]. As the number of GW events from CBCs increases, the lensing effect is expected to be detected on GWs soon [ng-18, li-18, oguri-18, xu-ezquiaga-holz-22, mukherjee-21, wierda-21].
The approaches to the observation of lensing of light and lensing of GWs are generally different, due to the nature of the detection methods. For light, lensing effects appear as either (1) multiple images of the same source, resolved through imaging (strong lensing [zwicky-37]); (2) transient increase of the source brightness following a Paczyński curve, when the source temporarily travels behind the lens [paczynski-86], seen through photometry (microlensing); (3) astrometric microlensing [paczynski-95]. For transient GWs detectable in ground-based detectors, the equivalent of strong lensing is the detection of multiple images of a static source in time domain [e.g., wang-96]. Given the transient nature of the detected compact binary merger (CBC) sources, the equivalent to microlensing by low mass lenses are wave optics distortions on static sources [nakamura-98, takahashi-03], rather than moving sources. The equivalent of the Paczyński curve is usually not considered for transient GWs from CBCs in ground-based detectors. Studies either consider a static source and varying frequency [e.g., liao-19, sun-fan-19, hou-20, biesiada-harikumar-21] or a constant frequency and a moving source, the latter of which has been analyzed for continuous GW sources [e.g., depaolis-01, guo-24, sudhagar-25, yang-25, li-25]. Joint studies have been done for space-based detectors [dorazio-loeb-20, yu-21]. To the best of the author’s knowledge, no previous studies have been done for a frequency-dependent moving source in ground-based detectors.
@@
The observation of the lensing curve can give us information of the parameters of the lensing system. In microlensing of light, the duration of the curve (twice the Einstein radius crossing timescale $t_{\rm E}$) usually has a degeneracy between the lens mass, a combination of the distances, and the relative velocity of the lens and the source [e.g., narayan-bartelmann-96, lee-17, gould-19]. The degeneracy may be broken by measuring the parallax [refsdal-66, gould-94b, gould-00], simultaneous parameters [lee-17, gould-19, for a review], finite-source effects [gould-94a, nemiroff-94], or relativistic effects [rahvar-20]. By contrast, when the source is orbiting a massive black hole (MBH) that acts as the lens (self-lensing), the Einstein radius crossing timescale is only dependent on the orbital distance $d_{\rm LS}$ from the source to the lens [rahvar-11]. For both light and GWs, the breaking of the Einstein crossing timescale degeneracy thus provides an independent way to measure the orbital distance $d_{\rm LS}$ of the source.
@@ II Lensing curve on a chirping signal
Figure 1: Schema of the lensing configuration considered in this study. Top: The GW source (S) is a compact binary orbiting a MBH at a distance $d_{\rm LS}$. The MBH acts as a gravitational lens (L), affecting the GW signal, with arrives distorted at the observer (O) when S, L and O are close to alignment. Bottom: View from the observer, where the source travels behind the lens. The lensing effect is dominant behind the cross-section of radius $R_{\rm E}$ and, as we will see in Sec. IV, the amplitude of the signal is modulated by the transmission factor ($F$ in Eq. (3)) shown by the oscillating pattern here. The source can have an offset $y_{0}R_{\rm E}$ with respect to the perfect alignment (dashed line going through the center of the lens).
The source travels behind the lensing cross-section, generally [^1] Following classical microlensing conventions, we are taking a maximum source position of $y_{\rm max}=1$ as the cross-section threshold. For GWs, the detectability thresholds for $y$ can be different (usually $\gtrsim 1$) than for electromagnetic lensing, and in that case the cross-section is $\pi R_{\rm E}^{2}y_{\rm max}^{2}$. given by $\pi R_{\rm E}^{2}$, where $R_{\rm E}=\sqrt{2R_{\rm S}d_{\rm LS}d_{\rm L}/d_{\rm S}}$ is the Einstein radius of the lens, which for self-lensing is simplified to $R_{\rm E}\simeq\sqrt{2R_{\rm S}d_{\rm LS}}$  [^2] Throughout this work, due to self-lensing, the distance $d_{\rm L}$ from the observer to the lens can be taken as practically equal to the distance $d_{\rm S}$ from the observer to the source., where $d_{\rm L}$ and $d_{\rm S}$ are the distance from the observer to the lens and to the source respectively. $R_{\rm S}=2GM_{\rm MBH}/c^{2}$ is the Schwarzschild radius of the MBH of mass $M_{\rm MBH}$, where $G$ is the gravitational constant and $c$ is the speed of light in vacuum. The source is generally not perfectly aligned (i.e., not crossing exactly behind the lens), it has a physical offset $\eta_{0}\simeq y_{0}R_{\rm E}$ where $y_{0}$ is the dimensionless source position (“impact parameter”, sometimes known as $u_{0}$) of closest approach to the lens at time $t_{0}$. As the source moves, the source position $y$ with respect to the lens changes over time ($t$),
@@ II.1 The crossing timescale depends directly on the orbital distance
Compared to a general source —where $t_{\rm E}$ depends on the combination of the distances $d_{\rm LS}d_{\rm L}/d_{\rm S}$, on the relative velocity between lens and source, and on the mass of the lens—, this result is not degenerate and provides a direct measure of the distance $d_{\rm LS}$, as we analyze in the following.
@@ II.2 Lensing signature on the GW waveform
A lensed CBC signal is shown in Fig. 2. To interpret it, we start chronologically with the source far from alignment, when the first image looks unlensed ($\sqrt{\mu_{1}}\rightarrow 1$), the second image is very faint ($\sqrt{\mu_{2}}\rightarrow 0$) and the time delay between them is large. As the source moves behind the lens, both images are magnified and the time delay shortens. At $t_{0}$, the images arrive simultaneously and at their maximum joint amplification, corresponding to the peak. As the source moves away, the situation is reversed back to the same as in the beginning, in this case the first image gets demagnified until it is too faint ($\sqrt{\mu_{1}}\rightarrow 0$) and the second image looks unlensed, $\sqrt{\mu_{2}}\rightarrow 1$.
@@
Figure 2: Representation of a GW signal CBC “chirp”, through the strain $h$ (relative deformation of space-time) as a function of time $t$. The amplitude of the Paczyński-like curve on the inspiral phase of the signal can be comparable to (or even larger than) the final merger. The parameters used here are a source “chirp mass” $M_{\rm chirp}=1\,{\rm M}_{\odot}$, a distance from the observer $d_{\rm L}=1\,{\rm Gpc}$ (just affecting the overall amplitude of the strain), $y_{0}=0.1$ and $d_{\rm LS}\simeq 3\times 10^{10}\,{\rm m}$. In self-lensing, the width of the lensing curve ($2t_{\rm E}=4d_{\rm LS}/c$) is independent on $M_{\rm MBH}$ [Eq. (2)]. Instead, as we will see in Sec. IV, $M_{\rm MBH}$ can be obtained from the interference pattern modulations on the signal (Fig. 3). For comparison purposes, the lensed signal is represented in the time reference of the arrival time of the first image, $t_{1}$. For simplicity, we represent the inspiral part of the signal from $f\gtrsim 12\,{\rm Hz}$, truncated at merger.
@@
Although the source waveform $h_{\rm UL}(t)$ is chosen to be a CBC “chirp”, with frequency dependence $f(t)\propto M_{\rm chirp}^{-5/8},$ the method can be applied to any smoothly varying source. In Fig. 2, the oscillations are too narrow to distinguish the interference pattern or the phase. A zoom-in view is shown in Fig. 3: while the interference pattern modulations are distinguishable, the oscillations still have too many cycles to be resolved by eye. In this work, the amplitude already gives us enough information. However, the phase may give additional information about environmental phase shifts, which are not taken into account here and may be required in more detailed searches.
@@
Figure 3: Zoom-in view of Fig. 2, at the peak of the lensing curve at GW frequencies $f\simeq 15\,{\rm Hz}$. The interference pattern modulating the GW signal depends on the mass, (a) $M_{{\rm MBH},z}=10^{5}\,M_{\odot}$, (b) $M_{{\rm MBH},z}=10^{6}\,M_{\odot}$, and the distance $d_{\rm LS}\simeq 3\times 10^{10}\,{\rm m}$ (or equivalently $t_{\rm E}\simeq 200\,{\rm s}$). The modulation period far from the peak can be estimated with Eq. (15), for example here in the segment $t\in(400-420)\,{\rm s}$.
@@
In ground-based GW detectors, this lensing curve on the GW signal may appear as an excess power. If not taken into account, the full signal may be missed or misinterpreted. Although in most cases the lensing signature might not be confused with short bursts (glitches or other burst signals such as eccentric pulses/bursts [samsing-25-GWpulsar]) because $t_{\rm E}$ is generally longer than their duration, the lensing curve should be taken into account when treating multiple overlapping signals in next-generation detectors.
@@ III Obtaining the orbital distance
Observability for sources of light. The light-curve may be seen for long timescales, provided that the source is at a distance $d_{\rm LS}$ larger than the tidal disruption radius of the source $R_{\rm tidal}$. For typical $R_{\rm tidal}$ of stars around MBHs, we can expect $t_{\rm E}\gtrsim 3\,{\rm hours}$. Low-mass MBHs may be favored due to their potential to produce short timescales because $d_{\rm LS}>R_{\rm tidal}\propto M_{\rm MBH}^{1/3}$, which can be interesting for the MBH at the center of our Galaxy (expanded in the Discussion, VI) or intermediate-mass black holes (IMBHs) at the centers of star clusters.
@@
Such close distances may be reached in active galactic nuclei (AGN) disks, if the sources are in migration traps [mckernan-12, bellovary-16, secunda-19]. For example, fixing the distance at $d_{\rm LS}\sim 10\,R_{\rm S}$ [peng-21, leong-25], for $M_{\rm chirp}=1\,M_{\odot}$, a complete curve appears when $M_{\rm MBH}\lesssim 3\cdot 10^{6}\,M_{\odot}$.
@@ IV Obtaining the MBH mass
The lensed GW signal presents a modulation of the amplitude which can be seen in Fig. 3 (a zoom-in view of Fig. 2). The modulation is caused by the phase difference between the lensed images, which creates an interference pattern that can also be found in continuous GW signals [sudhagar-25]. The period of this modulation can be estimated in the GO approximation from Eq. (7), $T=t/(f\Delta t)$ [savastano-24], using $y\propto t/t_{\rm E}$. This period is only formally valid for the radial spacing (interference pattern in Fig. 1). Since in our case the source has an offset $y_{0}$, the period will asymptotically match the modulations far from the $t_{0}$ peak. Nevertheless, the point of this estimation is to show that the period is inversely proportional to the mass of the MBH,
@@
By observing the modulation period $T$, the width of the curve $2t_{\rm E}$, and the GW frequency $f$  [^3] The frequency is assumed to be directly measurable by the detector from the incoming GWs., we can obtain the mass of the MBH:
@@
For the astrophysical parameters we consider (typical MBH masses and LVK/ET/CE frequencies), most times these modulations are expected to last $\gtrsim{\rm ms}$. The interference pattern is thus potentially resolvable in time domain in current ground-based GW detectors.